OpenStudy (anonymous):

Multiply and simplify:( 10x^2/6y^3 ) (24y^2/35x^2 )

OpenStudy (campbell_st):

\[\frac{5x^2\times2}{6y^2 \times y} timesr \frac{6y^2 \times 4}{5x^2 \times 7}\] cancel out the common factors leaving \[\frac{ 4 \times2}{y \times 7} = \frac {8}{7y}\]

OpenStudy (anonymous):

okay thanks :)